a. Calculate the sample average using ungrouped and grouped methods.

b. Calculate the sample standard deviation using ungrouped and grouped method
Answer #1

(a-b)

For un-grouped data:

Following table shows the calculations:

X (x-mean)^2
10 17.0569
7 50.8369
12 4.5369
11 9.7969
9 26.3169
13 1.2769
14 0.0169
10 17.0569
6 66.0969
12 4.5369
17 8.2369
17 8.2369
13 1.2769
15 0.7569
18 14.9769
14 0.0169
15 0.7569
16 3.4969
15 0.7569
14 0.0169
9 26.3169
19 23.7169
15 0.7569
14 0.0169
15 0.7569
16 3.4969
15 0.7569
14 0.0169
13 1.2769
16 3.4969
17 8.2369
13 1.2769
14 0.0169
11 9.7969
12 4.5369
15 0.7569
16 3.4969
13 1.2769
16 3.4969
15 0.7569
18 14.9769
15 0.7569
13 1.2769
15 0.7569
14 0.0169
16 3.4969
13 1.2769
16 3.4969
15 0.7569
16 3.4969
20 34.4569
14 0.0169
10 17.0569
15 0.7569
13 1.2769
15 0.7569
12 4.5369
14 0.0169
16 3.4969
13 1.2769
16 3.4969
13 1.2769
14 0.0169
16 3.4969
14 0.0169
15 0.7569
16 3.4969
15 0.7569
16 3.4969
15 0.7569
Total 989 441.843
14.12857 2.530518077

Sample size: n=70

The sample mean is:

\bar{x}=\frac{\sum x}{n}=\frac{989}{70}=14.13

The sample standard deviation is

s=\sqrt{\frac{\sum \left ( x-\bar{x} \right )^{2}}{n-1}}=\sqrt{\frac{441.843}{70-1}}\approx 2.53

For grouped data:

Following table shows the calculations:

Classes
Lower limit Upper limit Mid Point, X Frequency, f fx f(x-mean)^2
6 8.1 7.05 2 14.1 97.7202
8.1 10.2 9.15 5 45.75 119.5605
10.2 12.3 11.25 6 67.5 46.7046
12.3 14.4 13.35 21 280.35 9.9981
14.4 16.5 15.45 29 448.05 57.6549
16.5 18.6 17.55 5 87.75 61.6005
18.6 20.7 19.65 2 39.3 62.9442
Total 70 982.8 456.183

The sample mean is

\bar{x}=\frac{\sum fx}{\sum f}=14.04

The sample standard deviation is

s=\sqrt{\frac{\sum f(x-\bar{x})^{2}}{\sum f -1}}=\sqrt{\frac{456.183}{69}}=2.57